OpenStudy (countrygirl1431):

Four angles of a pentagon have measures of 85°, 90°, 95°, and 110°. The smallest of all the exterior angles measures is: 5° 20° 70° 85°

OpenStudy (beaucieldemerveilles):

70 degrees

OpenStudy (anonymous):

70 dgres

OpenStudy (countrygirl1431):

could yall help me with a few more

OpenStudy (anonymous):

yes

OpenStudy (beaucieldemerveilles):

sure, i'd be glad to help

OpenStudy (anonymous):

The interior and exterior angle add up to 180 degrees; so far the smallest would be 180 - 110 = 70 degrees. But since there are 5 angles in a pentagon, we have to first make sure that the unknown angle of the pentagon is not larger than 110 (making it's supplement, the exterior angle, smaller than 70 degrees. Draw a pentagon on your scratch paper (it doesn't have to be perfect) and pick one of the vertices (the "corners" where the sides meet). If you draw a line from that vertex to all of the other vertices, you see that you've divided the pentagon into 3 triangles. Since each triangle has 180 degrees, the whole pentagon has angles that add up to 3 x 180 or 540 degrees. From the four angles you've been told (85 + 90 + 95 + 110 = 380) you can figure out that the fifth angle has to be 160 degrees (540 - 380 = 160). So the exterior angle that matches up with that angle is 180 - 160 = 20 degrees and choice B. is correct! Also please Beau no direct answers and gurl000! If you need anymore math help please don't be scared to msg me or tag me in a question! Thank you, Have a nice day.

OpenStudy (anonymous):

can yall at lease help me new here

OpenStudy (countrygirl1431):

The midpoint of the segment joining points (a, b) and (j, k) is: ( j - a, k - b) (j + a, k + b)

OpenStudy (anonymous):

ummm

OpenStudy (anonymous):

a+j, k+b

OpenStudy (countrygirl1431):

i thought so

OpenStudy (anonymous):

sketch it on paper, u nwill c

OpenStudy (countrygirl1431):

Ray bisects AOC and AOX = 42°. AOC = 42° 84° 21° 68°

OpenStudy (countrygirl1431):

ray OX

OpenStudy (anonymous):

would say 84

OpenStudy (countrygirl1431):

thanks

OpenStudy (anonymous):

but then its ray oa that bisects it

OpenStudy (anonymous):

yay 84

OpenStudy (anonymous):

if thats the case then its 42

OpenStudy (anonymous):

so 84 / 2

OpenStudy (anonymous):

for this type of exercises a drawning needs to be sketched to make it visible

OpenStudy (anonymous):

got another problem?